A helicopter blade starts to rotate from rest with a constant angular acceleration of \alpha=\:
α
=
0.76 radians/second2. How many revolutions will the blade make before it is rotating at 500 revolutions per minute (RPM)?
Answer #1

wf = 500 * 2 pi / 60 = 52.35 rad/s

wf2 - wi2 = 2 * alpha * theta

52.352 = (2 * 0.76 * n * 2 pi)

n = 52.352 / (2 * 0.76 * 2 pi)

n = 287 rev